Prove that a strictly decreasing function is always one-one.


1
Expert's answer
2021-04-14T05:44:49-0400

"Solution:
\\Proof~1:
\\For~ a ~strictly~ decreasing ~function~ if ~x_1<x_2 ~then ~ f(x_1)>f(x_2).
\\Here ~no ~ two ~ values ~ of ~ domain ~ can ~ have ~ same ~values ~ in ~ codomain.
\\ One ~may ~ be ~ less ~than ~ or ~ greater ~than ~ other ~but ~ not ~equal .So~ a ~strictly ~
\\ decreasing ~ function ~ is ~always one-one.

\\Proof ~ 2:
\\Given: f:R\rightarrow~R~ is ~strictly~ decreasing ~function.
\\To ~ prove: f(x)~ is ~ one-one.
\\proof: f~ is ~strictly ~ decreasing ~ implies: if~ x<y, ~then ~f(x)>f(y)
\\Let ~ us ~ assume ~ that~ f(a)=f(b)
\\If ~ a<b, then ~by ~ the ~ definition ~ of ~ strictly~ increasing ~f(a)>f(b).
\\Thus~ it ~ is ~ not ~possible ~that ~a<b~ when ~ f(a)=f(b).
\\If ~ b<a, then ~by ~ the ~ definition ~ of ~ strictly~ increasing ~f(b)>f(a).
\\Thus~ it ~ is ~ not ~possible ~that ~b<a~ when ~ f(a)=f(b).
\\Since ~ a<b ~ is ~ not ~true ~ and ~ since ~ b<a ~ is ~ not ~ true ,a ~and~ b ~then ~ have ~ to ~be ~\\ the ~ same~: 
\\~~~~~~~~~~~~~~~~~~~~~~~~~~a=b
\\By~ definition ~ of ~one-to-one~ function, ~we~ have ~ then ~ shown ~ that ~f~is ~
\\one-to-one ."